OpenStudy (anonymous):

9th grade math, exponential functions. ex: 25 to the (10x+8) is less than or equal to (1 over 125) to the (4-2x) power

OpenStudy (anonymous):

\[25^{10x-8}\ge \left(\begin{matrix}1 \\ 125\end{matrix}\right)^{4-2x}\]

OpenStudy (anonymous):

what the equation looks like ^^^